I have two problems i need help with. I need to solve these by using the quadratic equation.
1: 3x^2 - 2x = 15x - 0
2: x^2 - 3x - 23 = 5

Any help would be greatly appreciated

Answer by checkley71(8403)

3X~2-2X-15X=0 OR 3X~2-17X=0 OR X(3X-17) OR X=0 & 3X=17 OR X=17/3
X~2-3X-23-5=0 OR X~2-3X-28 OR (X-7)(X+4)=0 OR X=7 & X=-4